# Collision-Free Navigation of Wheeled Mobile Robots: An Integrated Path Planning and Tube-Following Control Approach

### v1

@@ Abstract
In this paper, an integrated path planning and tube-following control scheme is proposed for collision-free navigation of a wheeled mobile robot (WMR) in a compact convex workspace cluttered with sufficiently separated spherical obstacles. An analytical path planning algorithm is developed based on Bouligand’s tangent cones and Nagumo’s invariance theorem, which enables the WMR to navigate towards a designated goal location from almost all initial positions in the free space, without entering into augmented obstacle regions with safety margins. We further construct a virtual “safe tube” around the reference trajectory, ensuring that its radius does not exceed the size of the safety margin. Subsequently, a saturated adaptive controller is designed to achieve safe trajectory tracking in the presence of disturbances. It is shown that this tube-following controller guarantees that the WMR tracks the reference trajectory within the predefined tube, while achieving uniform ultimate boundedness of both the position tracking and parameter estimation errors. This indicates that the WMR will not collide with any obstacles along the way. Finally, we report simulation and experimental results to validate the effectiveness of the proposed method.
@@
Tube-following control, adaptive control, path planning, robot navigation, wheeled mobile robots.
@@ I Introduction
The path planning and control of wheeled mobile robots (WMRs) is an important problem in robotics due to the extensive real-world applications that these systems have, such as cargo transportation, exploration of hazardous environments, automated patrolling [1], to name a few cases. In practical applications, WMRs usually operate in obstacle-cluttered environments, which limits the implementation of traditional motion control methods, thus, motivates the development of collision-free navigation algorithms that can steer WMRs from an initial position to a target goal without colliding with any obstacles along the way [2]. Collision-free robot navigation has garnered significant attention from the robotics and control research communities [3]. Existing solutions to the collision-free navigation problem of robots can be primarily classified into two categories [4]: global (map-based) methods and local (reactive) methods. The former relies on global knowledge of the environment, requiring a priori information about the obstacles (e.g., position and shape). In contrast, the latter utilizes only local knowledge of obstacles in the immediate vicinity of the robot, obtained from on-board sensors. Among the global methods, a simple and computationally efficient solution is the artificial potential field (APF)-based approach [5], which uses a field of virtual potential forces to push the robot towards a target position and pull it away from obstacles; Representative APF-based methods can be found in [6, 7, 8] and references therein. These kind of methods often reach local minima, which hinders achieving global convergence to a designated goal (especially in topologically complex settings). Although navigation functions developed in [9] can help to deal with this problem, they come with extra drawbacks, such as the need for unknown tuning of parameters. Other global methods for collision-free navigation are e.g., heuristic approaches like $A^{\star}$ [10, 11], rapidly exploring random tree (RRT) [12], genetic algorithms [13], etc. As most of them are search-based solutions, their performance can be significantly affected by the scale of the problem [14]. Optimization-based methods [15], while providing alternatives, typically rely on numerical solutions of the constrained optimization problems that result on heavily computational costs. Both the search-based and optimization-based methods are inherently open loop, which makes them vulnerable to disturbances and sensor noise. Thus, the actual robot trajectory may deviates from the desired path, hence, risking collisions with the environment. Local methods that offer reactive solutions for collision-free robot navigation are highly desirable in practice. Particularly in autonomous exploration applications, where WMRs often have limited access to global information and must rely on local sensory data to detect obstacles and perform navigate. Bug algorithms that evolved from maze solving algorithms are one of the simplest reaction navigation approaches for mobile robots in planar environments [16, 17]. Lionis et al. [18] proposed locally computable navigation functions for robotic navigation in unknown sphere worlds. Filippidis and Kyriakopoulos [19] introduced adjustable navigation functions that are capable of autonomously tuning parameters when new obstacles are discovered. Paternain and Ribeiro [20] provided a stochastic extension of these navigation functions with provable convergence properties. However, one caveat is that these reactive function-based methods also suffer from local minima [4]. Arslan and Koditschek [21] proposed a sensor-based reactive method based on a robot-centric spatial decomposition for collision-free navigation with convex obstacles. Huber et al. [22] adopted contraction-based dynamical systems theory to achieve a closed-form solution to the avoidance of moving obstacles. Recently, reinforcement learning methods were used in [23], however, the learning process requires a substantial amount of interaction with the real world. Note that although most of the aforementioned reactive approaches are closed-form, they only consider nominal (holonomic or non-holonomic) robot kinematics, without taking into account disturbances/uncertainties that may arise from external forces, modeling errors, inaccurate controls, sensor noise, data dropout, etc. These unknown actions/terms inevitably degrade the system’s performance, potentially leading to deviations of the desired trajectory, thus, jeopardizing the robot’s safety. In this paper, we study the collision-free navigation problem of differential-drive WMRs operating in a compact convex workspace cluttered with multiple sufficiently separated spherical obstacles. To overcome the non-holonomic constraints inherent in this type of systems, an off-center point is selected as the virtual control point, enabling to transform the original non-holonomic kinematics into a fully-actuated two degrees-of-freedom system [24, 25]. Building upon this model, we propose a novel robot navigation scheme that integrates a path planner and a tube-following controller. By “tube-following” it is meant that the tracking trajectory of the WMR should remain within a virtual safe tube around the reference trajectory to ensure its motion safety. The proposed method can achieve collision-free navigation in cluttered environments from almost all initial configurations in the robot’s free space despite the presence of disturbances. To the best of the authors’ knowledge, this is the first time an integrated path planning and tube-following control scheme is proposed for safe robot navigation in cluttered environments. The effectiveness of the proposed method is validated through numerical simulations and experiments. Specifically, the original contributions of the paper are two-fold:
@@
    Inspired by [4], a safe path planner is presented based on Bouligand’s tangent cones [26] and Nagumo’s invariance theorem [27], yielding a continuous saturated vector field that can guide the WMR from almost all initial positions in the free space towards a designated goal, without entering into the augmented obstacle regions with safety margins. In essence, the proposed planning algorithm is sensor-based, and the vector field can be generated in real-time using only the local sensing information (i.e., the distance between the WMR and the obstacle boundaries).
@@
    We propose an adaptive tube-following control method that defines a virtual safe tube around the reference trajectory with its radius no larger than the size of the safety margin. It is shown that the derived controller ensures the tracking of the reference trajectory within the predefined safe tube, while achieving uniform ultimate boundedness of both the trajectory tracking and parameter estimation errors, despite the presence of disturbances. By design, the chosen tube radius ensures that the WMR’s trajectory never enters into the actual obstacle regions, thus achieving the collision-free robot navigation in the perturbed conditions.
@@ II-A Kinematics of wheeled mobile robots
where $\mathbf{u}:=[v,\omega]^{\top}\in\mathbb{R}^{2}$ is the the control input vector with $v$ and $\omega$ being the linear and angular velocities of the WMR, respectively, and $\mathbf{u}_{d}\in\mathbb{R}^{2}$ is the matched disturbance that may arise from control errors, sensor noise, etc.
@@
To facilitate the subsequent design and analysis, we select an off-center point $\bm{P}$ as the virtual control point, which locates at a distance $\ell\neq 0$ away from $\bar{\bm{P}}$ along the longitudinal axis of the WMR [24, 25], as shown in Fig. 1. Denote by $\mathbf{x}:=[x,y]^{\top}\in\mathbb{R}^{2}$ and $\theta$ the position and heading angle of $\bm{P}$, respectively. We have the following change of coordinates:
@@
where $\mathbf{R}(\theta)\in\mathbb{R}^{2\times 2}$ is defined as
$\mathbf{R}(\theta):=\left[\begin{matrix}\cos\theta&-\ell\sin\theta\\ \sin\theta&\ell\cos\theta\\ \end{matrix}\right],$ (4)
which is full rank. By setting $|\ell|\leq 1$, it follows that $\|\mathbf{R}(\theta)\|=\max\{1,|\ell|\}=1$ and $\|\mathbf{R}^{-1}(\theta)\|=\max\{1,1/|\ell|\}=1/\ell$.
@@
The full rank of $\mathbf{R}(\theta)$ enables us to freely control the position of the WMR (specifically, the virtual control point $\bm{P}$) without being restricted by nonholonomic constraints.
@@ II-B Operating environment
Consider a WMR operating inside a closed compact convex workspace $\mathcal{W}^{\ast}\subset\mathbb{R}^{2}$, punctured by a set of $n\in\mathbb{N}$ obstacles $\mathcal{O}_{i}^{\ast}$, $i\in\mathbb{I}:=\{1,2,...,n\}$, which are represented by open balls with centers $\mathbf{c}_{i}\in\mathbb{R}^{2}$ and radii $r_{i}>0$. A circumscribed circle centered at $\bm{P}$ with radius $r>0$ is constructed, which is the smallest circle enclosing the WMR (see Fig. 1). To ensure that the WMR can navigate freely between any of the obstacles in $\mathcal{W}^{\ast}$, we make the following assumption:
@@
For ease of design, the WMR is considered as a point (i.e., the virtual control point $\bm{P}$), by transferring the volume of the circumscribed circle to the other workspace entities. Then, for the point $\bm{P}$, the workspace is defined as
$\mathcal{W}:=\{\mathbf{x}\in\mathbb{R}^{2}\mid\text{d}_{\complement\mathcal{W}^{\ast}}(\mathbf{x})\geq r\},$ (7)
and the obstacle regions are defined as the following pairwise disjoint spherical subsets of $\mathbb{R}^{2}$:
$\mathcal{O}_{i}:=\{\mathbf{x}\in\mathbb{R}^{2}\mid\beta_{i}(\mathbf{x})<0\},~{}i\in\mathbb{I},$ (8)
where $\beta_{i}(\mathbf{x}):=\|\mathbf{x}-\mathbf{c}_{i}\|-(r+r_{i})$. Thus, the free space of $\bm{P}$ is given by the closed set $\mathcal{X}:=\mathcal{W}\setminus\mathcal{O}$ with $\mathcal{O}:=\cup_{i=1}^{n}\mathcal{O}_{i}$. To enhance the safety, a safety margin of size $0<\epsilon<h$ is introduced (see the light blue regions in Fig. 2), which results in an eroded workspace $\mathcal{W}^{\epsilon}:=\{\mathbf{x}\in\mathbb{R}^{2}\mid\text{d}_{\complement\mathcal{W}^{\ast}}(\mathbf{x})\geq r+\epsilon\}$ and $n$ augmented obstacle regions $\mathcal{O}_{i}^{\epsilon}:=\{\mathbf{x}\in\mathbb{R}^{2}\mid\beta_{i}(\mathbf{x})<\epsilon\}$, $i\in\mathbb{I}$. By doing so, the free space of $\bm{P}$ reduces to $\mathcal{X}_{\epsilon}:=\mathcal{W}^{\epsilon}\setminus\mathcal{O}^{\epsilon}$ with $\mathcal{O}^{\epsilon}:=\cup_{i=1}^{n}\mathcal{O}_{i}^{\epsilon}$.
Figure 2: Schematic diagram of the operating environment. The dark gray balls denote the obstacles $\mathcal{O}_{i}^{\ast}$, $i\in\mathbb{I}$, the light gray regions denote the augmented workspace boundary and obstacles, whereas the light blue regions denote the safety margin.
@@ II-C Problem formulation
The collision-free navigation problem for the WMR is now formulated as follows:
@@
[…]
@@ III Main Results
To solve Problem 1, we propose a closed-form robot navigation scheme that integrates a path planner and a tube-following controller. The block diagram of this scheme is depicted in Fig. 3. The path planner explicitly considers the safety margin and generates a collision-free reference trajectory $\mathbf{x}_{d}$ in the free space $\mathcal{X}{\epsilon}$, using tangent cones and the well-known Nagumo’s invariance theorem. In the control module, an adaptive tube-following controller is designed to track the planned trajectory in the presence of unknown disturbances, while ensuring that the position $\mathbf{x}$ of $\bm{P}$ stays within a predefined tube around the reference trajectory $\mathbf{x}_{d}$. Under this framework, if the radius of the tube is set less than the width $\epsilon$ of the safety margin, then the actual motion trajectory of $\mathbf{x}$ will remain within the set $\mathcal{X}$, thus achieving safe robot motions.
Figure 3: Block diagram of the collision-free navigation scheme.
@@ III-A Preliminaries
The definition of tangent cones and Nagumo’s theorem that are necessary for collision-free path planning are given.
@@
$\mathbf{T}_{\mathcal{F}}(\mathbf{x}):=\left\{\mathbf{z}\in\mathbb{R}^{n}\mid\mathop{\lim\inf}\limits_{\tau\to 0^{+}}\dfrac{{\rm d}_{\mathcal{F}}(\mathbf{x}+\tau\mathbf{z})}{\tau}=0\right\}.$
@@
Consider the system $\dot{\mathbf{x}}(t)=\mathbf{f}(\mathbf{x}(t))$ and assume that it admits a unique solution in forward time for each initial condition $\mathbf{x}(0)$ in an open set $\mathcal{O}$. Then, the closed set $\mathcal{F}\subset\mathcal{O}$ is forward invariant iff
$\mathbf{f}(\mathbf{x})\in\mathbf{T}_{\mathcal{F}}(\mathbf{x}),~{}\forall\mathbf{x}\in\mathcal{F}.$ (9)
An intuitive interpretation of Nagumo’s theorem is that if the vector field $\mathbf{f}(\mathbf{x})$ points inside or is tangent to the set $\mathcal{F}$ at each point $\mathbf{x}\in\mathcal{F}$, then $\mathcal{F}$ is forward invariant. As stated earlier, $\mathbf{T}_{\mathcal{F}}(\mathbf{x})$ is non-trivial only on $\partial\mathcal{F}$, thus it is only necessary to check (9) for the boundary points.
Lemma 1.
Given any constant $b>0$, the following inequalities hold for all $0\leq q<b$:
$\frac{q}{b}\leq\ln\dfrac{b}{b-q}\leq\dfrac{q}{b-q}.$
Proof. Define two functions $h_{1}(q):=q/(b-q)-\ln(b/(b-q))$ and $h_{2}(q):=\ln(b/(b-q))-q/b$. After simple calculations, it follows that
$\dfrac{\partial h_{1}(q)}{\partial q}=\dfrac{q}{(b-q)^{2}}\geq 0,~{}\dfrac{\partial h_{2}(q)}{\partial q}=\dfrac{q}{b(b-q)}\geq 0,$
[…]
III-B Path planing based on tangent cones
Based on Bouligand’s tangent cones and Nagumo’s theorem, a safe path planning approach is proposed, which can guide the virtual control point $\bm{P}$ towards the goal $\mathbf{x}^{\ast}$ (motion-to-goal) from almost [^1] The free space $\mathcal{X}_{\epsilon}$ is a non-contractible space, for which the topological obstruction precludes the possibility of any continuous vector fields achieving global asymptotic stability. The basin of attraction of the desired equilibrium must generally exclude at least a set of measure zero. all initial conditions $\mathbf{x}(0)\in\mathcal{X}_{\epsilon}$, while ensuring forward invariance of the free space $\mathcal{X}_{\epsilon}$ (collision avoidance). We neglect the disturbance $\mathbf{u}_{d}$ in (3) and consider a control law of the following form
$\mathbf{u}=\mathbf{R}^{-1}(\theta)\bm{\tau},$ (10)
where $\bm{\tau}\in\mathbb{R}^{2}$ is a virtual control law to be designed. Then, substituting (10) into (3), one gets

### v2

# InPTC: Integrated Planning and Tube-Following Control for Prescribed-Time Collision-Free Navigation of Wheeled Mobile Robots

@@ Abstract
In this article, we propose a novel approach, called InPTC (Integrated Planning and Tube-Following Control), for prescribed-time collision-free navigation of wheeled mobile robots in a compact convex workspace cluttered with static, sufficiently separated, and convex obstacles. A path planner with prescribed-time convergence is presented based upon Bouligand’s tangent cones and time scale transformation (TST) techniques, yielding a continuous vector field that can guide the robot from almost all initial positions in the free space to the designated goal at a prescribed time, while avoiding entering the obstacle regions augmented with safety margin. By leveraging barrier functions and TST, we further derive a tube-following controller to achieve robot trajectory tracking within a prescribed time less than the planner’s settling time. This controller ensures the robot moves inside a predefined “safe tube” around the reference trajectory, where the tube radius is set to be less than the safety margin. Consequently, the robot will reach the goal location within a prescribed time while avoiding collision with any obstacles along the way. The proposed InPTC is implemented on a Mona robot operating in an arena cluttered with obstacles of various shapes. Experimental results demonstrate that InPTC not only generates smooth collision-free reference trajectories that converge to the goal location at the preassigned time of $250\,\rm s$ (i.e., the required task completion time), but also achieves tube-following trajectory tracking with tracking accuracy higher than $0.01\rm m$ after the preassigned time of $150\,\rm s$. This enables the robot to accomplish the navigation task within the required time of $250\,\rm s$.
@@
Prescribed-time control, path planning, trajectory tracking, collision avoidance, wheeled mobile robots.
@@ I Introduction
The motion control of wheeled mobile robots (WMRs) is a benchmark problem in robotics due to its key role in extensive real-world applications, such as cargo transportation, automated patrolling, and exploration of hazardous environments [1]. In these applications, WMRs typically operate in obstacle-cluttered environments, which motivates the development of safe navigation algorithms that can steer WMRs from an initial position to a desired goal location without colliding with any obstacles along the way [2]. Collision-free navigation of WMRs have garnered significant attention from the robotics and control research communities [3].
@@
Existing solutions to the robot navigation problem can be primarily classified into two categories [4]: global (map-based) methods and local (reactive) methods. The former requires a priori information of environments (e.g., position and shape of obstacles), while the latter utilizes only local knowledge of obstacles obtained from on-board sensors. Among the global methods, a computationally efficient solution is the artificial potential field (APF)-based approach [5], which uses a field of potential forces to push the robot towards a goal position and pull it away from obstacles [6, 7, 8]. Control barrier function (CBF) combined with quadratic programming (CBF-QP) also provides an effective strategy for developing controllers that can both stabilize a system and ensure safety, making it particularly useful for autonomous vehicles operating in obstacle-cluttered environments [9, 10, 11]. Note that both the APF- and CBF-based methods often reach local minima, which hinders achieving global convergence to the designated goal, especially in topologically complex settings. Although navigation functions presented in [12] can overcome this issue, they require unknown tuning of parameters. Other global methods include heuristic approaches such as $A^{\star}$ [13], rapidly exploring random tree (RRT) [14], and genetic algorithms [15]. Since these methods are search-based solutions, their performance is directly affected by the problem scale [16]. Optimization-based methods [17] provide alternatives, but typically require numerically solving constrained optimization problems, which results in high computational costs. Local methods that offer reactive solutions for collision-free robot motions are highly desirable, particularly in autonomous exploration applications, where robots often have limited access to environment information and must rely on local sensory data to detect obstacles. Bug algorithm is a simple reaction planning approaches for mobile robots [18, 19]. Arslan and Koditschek [20] proposed a reactive method based on a robot-centric spatial decomposition for collision-free navigation. Huber et al. [21] adopted contraction-based dynamical systems theory to achieve dynamic obstacle avoidance. Berkane et al. [4] proposed a sensor-based navigation algorithm that employs Nagumo’s invariance theorem to guarantee robot safety by projecting the nominal velocity onto safe velocity cones (e.g., Bouligand’s tangent cones) when the robot is close to obstacle boundaries. Although most reactive approaches offer closed-form solutions, they typically consider only the nominal robot kinematics, without accounting for disturbances that may arise, for example, from sensor noise, imperfect wheel alignment, and motor or actuator asymmetry. These unmodeled factors inevitably degrade system performance, potentially leading to deviations of the desired trajectory and compromising the robot’s safety. Furthermore, the aforementioned methods can only guarantee that WMRs reach the desired goal asymptotically, which is undesirable for time-critical missions.
There has been extensive research on the trajectory tracking control of WMRs. Zhai and Song [22] combined the adaptive and sliding mode control techniques to address the trajectory tracking problem of WMRs. Zheng et al. [23] presented an adaptive sliding mode control method for trajectory tracking of WMRs, and introduced a barrier function-like control gain to prevent input saturation. Kim and Singh[24] proposed a differential-flatness-based feedback controller to achieve the tracking of prescribed point-to-point trajectories. Recently, Zhou et al.[25] proposed a homogeneous control strategy to solve the trajectory tracking problem for perturbed unicycle mobile robots, where the convergence rate can be tuned by selecting a proper homogeneity degree. While these methods are effective in addressing stability and steady-state performance, they often do not explicitly consider the critical requirement of maintaining the WMR’s actual motion within a predefined safe tube around the reference trajectory. This actually imposes constraints on the tracking errors. Furthermore, most of these methods can only achieve asymptotic position tracking, and they typically lack the capability to preassign the convergence time according to the task requirements.
This paper examines the planning and control problems for prescribed-time collision-free navigation of WMRs in a convex workspace cluttered with static, sufficiently separated, and convex obstacles. By selecting an off-axis point as the virtual control point, the non-holonomic kinematics of WMRs is transformed into a fully-actuated system [26, Section 11.6.1]). Building upon this model, we propose a novel approach, called InPTC (Integrated Planning and Tube-Following Control), to achieve collision-free navigation of WMRs within a finite time that can be preassigned according to the task requirement, even in the presence of disturbances. Numerical simulations and experiments illustrate the effectiveness of the proposed InPTC. The main contributions of this paper are twofold:
@@
    A reactive path planner with prescribed-time convergence is developed based on Bouligand’s tangent cones [4] and time scale transformation [27]. This planner yields a continuous vector field that can guide the WMR from almost all initial positions in the free space to reach the goal at a prescribed finite time, while avoiding entering the obstacle regions augmented by safety margins.
@@
    By leveraging time scale transformation and barrier functions, a prescribed-time tube-following controller is derived for reference trajectory tracking. This controller guarantees that the tracking error converges to a small residual set within a prescribed finite time, whilst the WMR’s position remains within a predefined safe tube around the reference trajectory, despite the presence of disturbances.
The remainder of the paper is organized as follows: In Section II, we describe the kinematics and operating environment of WMRs, and formulate the prescribed-time safe navigation problem. In Section III, an integrated planning and control framework is proposed to address this problem. The simulation and experimental results are presented in Sections IV and V, respectively. Finally, Section VI provides concluding remarks and discusses future work.
@@ II-A Kinematics of wheeled mobile robots
where $\mathbf{u}:=[v,\omega]^{\top}\in\mathbb{R}^{2}$ is the the control input vector with $v$ and $\omega$ being the linear and angular velocities of the WMR, respectively, and $\mathbf{u}_{d}\in\mathbb{R}^{2}$ is the matched disturbance that may arise from sensor noise, imperfect wheel alignment, and motor or actuator asymmetry.
@@
To facilitate the subsequent design and analysis, we select an off-axis point $\bm{P}$ as the virtual control point, which locates at a distance $\ell\neq 0$ away from $\bar{\bm{P}}$ along the longitudinal axis of the WMR [26], as shown in Fig. 1. Denote by $\mathbf{x}:=[x,y]^{\top}\in\mathbb{R}^{2}$ and $\theta$ the position and heading angle of $\bm{P}$, respectively. We have the following change of coordinates:
@@
where $\mathbf{R}(\theta):=[\cos\theta,-\ell\sin\theta;\sin\theta,\ell\cos\theta]$ is full rank. This enables us to freely steer the WMR’s position regardless of nonholonomic constraints. By setting $|\ell|\leq 1$, it follows that $\|\mathbf{R}(\theta)\|=\max\{1,|\ell|\}=1$.
@@
The kinematics given by (1) represents a general kinematic model for unicycle mobile robots, such as differential drive robots and synchro drive robots. Additionally, following the input/output linearization outlined in [26, Section 11.6.1], the nonholonomic kinematics (1) is transformed into a full-actuated system (3), via a change of coordinates given by (2). This transformed model can be rewritten in the form of a single integrator system, which has the same structure as the kinematics of omnidirectional robots with holonomic constraints. Consequently, the proposed method building upon (3) is applicable for a broad class of WMRs.
@@ II-B Operating environment
Consider a WMR operating inside a closed compact convex workspace $\mathcal{W}^{\ast}\subset\mathbb{R}^{2}$, punctured by a set of $n$ static obstacles $\mathcal{O}_{i}^{\ast}$, $i\in\mathbb{I}:=\{1,2,...,n\}$, which are represented by open balls with centers $\mathbf{c}_{i}\in\mathbb{R}^{2}$ and radii $r_{i}>0$. A circumscribed circle centered at $\bm{P}$ with radius $r>0$ is constructed, which is the smallest circle enclosing the WMR (see Fig. 1). To ensure that the WMR can navigate freely between any of the obstacles in $\mathcal{W}^{\ast}$, we make the following assumption [4, 20]:
@@
For ease of design, the WMR is considered as a point (i.e., the off-axis point $\bm{P}$), by transferring the volume of the circumscribed circle to the other workspace entities. For the point $\bm{P}$, the workspace is $\mathcal{W}:=\{\mathbf{x}\in\mathbb{R}^{2}\mid\text{d}_{\complement\mathcal{W}^{\ast}}(\mathbf{x})\geq r\}$, and the obstacle regions are $\mathbb{R}^{2}$: $\mathcal{O}_{i}:=\{\mathbf{x}\in\mathbb{R}^{2}\mid\beta_{i}(\mathbf{x})<0\},~{}i\in\mathbb{I}$, where $\beta_{i}(\mathbf{x}):=\|\mathbf{x}-\mathbf{c}_{i}\|-(r+r_{i})$. Thus, the free space of $\bm{P}$ is given by the closed set $\mathcal{X}:=\mathcal{W}\setminus\mathcal{O}$ with $\mathcal{O}:=\cup_{i=1}^{n}\mathcal{O}_{i}$. To enhance the safety, a safety margin of size $0<\epsilon<h$ is introduced (see the light blue regions in Fig. 2), which results in an eroded workspace $\mathcal{W}^{\epsilon}:=\{\mathbf{x}\in\mathbb{R}^{2}\mid\text{d}_{\complement\mathcal{W}^{\ast}}(\mathbf{x})\geq r+\epsilon\}$ and $n$ augmented obstacle regions $\mathcal{O}_{i}^{\epsilon}:=\{\mathbf{x}\in\mathbb{R}^{2}\mid\beta_{i}(\mathbf{x})<\epsilon\}$, $i\in\mathbb{I}$. Then, the free space of $\bm{P}$ reduces to $\mathcal{X}_{\epsilon}:=\mathcal{W}^{\epsilon}\setminus\mathcal{O}^{\epsilon}$ with $\mathcal{O}^{\epsilon}:=\cup_{i=1}^{n}\mathcal{O}_{i}^{\epsilon}$.
Figure 2: Schematic diagram of the operating environment, where the dark gray balls denote the actual obstacles $\mathcal{O}_{i}^{\ast}$, the light gray regions denote the augmented obstacle regions $\mathcal{O}_{i}$, and the light blue regions denote the safety margin. In addition, the dashed circles are the influence regions of obstacles, whereas the pink and yellow lines denote the discontinuous trajectory and its continuous counterpart, respectively.
@@ II-C Problem formulation
The robot navigation problem is formulated as follows:
@@
Consider the WMR kinematics described by (3) under Assumptions 1 and 2. The objective is to derive a control law $\mathbf{u}$ that derives the virtual control point $\bm{P}$ from an initial position $\mathbf{x}(0)\in\mathcal{X}_{\epsilon}$ to a designated goal $\mathbf{x}^{\ast}\in\text{int}(\mathcal{X}_{\epsilon})$, without colliding with any obstacles along the route.
@@ III Main Results
In this section, an approach, called InPTC, is proposed to solve Problem 1. The block diagram of this scheme is shown in Fig. 3, where the planner module generates a collision-free reference trajectory $\mathbf{x}_{d}$ with prescribed-time convergence. In the control module, a tube-following controller is derived to achieve prescribed-time trajectory tracking within a safe tube around $\mathbf{x}_{d}$ (with radius less than the safety margin). Under this framework, $\mathbf{x}$ will converge to the goal $\mathbf{x}^{\ast}$ within a prescribed time, without colliding with any obstacles along the way, thus achieving prescribed-time safe robot navigation.
Figure 3: Block diagram of the proposed InPTC scheme.
@@ III-A Preliminaries
$\mathbf{T}_{\mathcal{F}}(\mathbf{x}):=\left\{\mathbf{z}\in\mathbb{R}^{n}\bigg{|}\mathop{\lim\inf}\limits_{\tau\to 0^{+}}\dfrac{{\rm d}_{\mathcal{F}}(\mathbf{x}+\tau\mathbf{z})}{\tau}=0\right\}.$
@@
Consider the system $\dot{\mathbf{x}}(t)=\mathbf{f}(\mathbf{x}(t))$, which admits a unique solution in forward time for each initial condition $\mathbf{x}(0)$ in an open set $\mathcal{O}$. The closed set $\mathcal{F}\subset\mathcal{O}$ is forward invariant iff $\mathbf{f}(\mathbf{x})\in\mathbf{T}_{\mathcal{F}}(\mathbf{x})$, $\forall\mathbf{x}\in\mathcal{F}$.
Definition 2 (Time scale transformation function, TSTF [27]).
A smooth function $\eta:[0,\infty)\to[0,T)$ is called a TSTF if 1) it is strictly increasing; 2) it is s.t. $\eta(0)=0$ and $\eta^{\prime}(0)=1$; 3) it is s.t. $\lim_{s\to\infty}\eta(s)=T$ and $\lim_{s\to\infty}\eta^{\prime}(s)=0$.
A TSTF squeezes the infinite-time interval $s\in[0,\infty)$ into a prescribed finite time interval $t:=\eta(s)\in[0,T)$, which plays a key role for achieving prescribed-time planning and control. In this work, $\eta(s)$ is of the form
$\eta(s)=T(1-e^{-\frac{s}{T}}).$ (6)
III-B Prescribed-time path planning
We neglect the disturbance $\mathbf{u}_{d}$ in (3) and consider a control law $\mathbf{u}=\mathbf{R}^{-1}(\theta)\bm{\tau}$, where $\bm{\tau}\in\mathbb{R}^{2}$ is a virtual control law to be designed. Then, substituting it into (3), one gets